OpenStudy (anonymous):

The perimeter of a rectangle is twice the sum of its length and its width. The perimeter is 16 meters and its length is 2 meters more than twice its width. What is its length? @misz.lopez115 @hartnn

hartnn (hartnn):

P = 16 = 2 (l+w) l = 2 + 2w solve :)

OpenStudy (anonymous):

So which equation do I use Hartnn?

OpenStudy (anonymous):

Confused ._.

OpenStudy (anonymous):

p=16 which then will qual to 2 which looks like (l=w) length plus width :) Then you do length okay? 2+2 times the width which will give you the answer :)

hartnn (hartnn):

both! 16 = 2 (l+w), so 8 = l +w , so w= 8-l plug this in l = 2 + 2w

OpenStudy (anonymous):

Ohhhh okay I get it now xD. Thanks.. Feel dumb for asking this now. Thanks to both of you, I got it now :) thanks @hartnn and @misz.lopez115

OpenStudy (joftheworld):

I still don't get it D: